OpenStudy (anonymous):

what is the x-intercept of y=-2x^2-32x-126

OpenStudy (anonymous):

y = 2x² - 8x + 6 y-intercept is (0,6) For x-intercepts... 0 = 2(x² - 4x + 3) 0 = 2(x - 3)(x - 1) x = {3,1}